# Extracting Probabilistic Knowledge from Large Language Models for Bayesian Network Parameterization

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated potential as factual knowledge bases; however, their capability to generate probabilistic knowledge about real-world events remains understudied. This paper investigates using probabilistic knowledge inherent in LLMs to derive probability estimates for statements concerning events and their interrelationships captured via a Bayesian Network (BN). Using LLMs in this context allows for the parameterization of BNs, enabling probabilistic modeling within specific domains. Experiments on eighty publicly available Bayesian Networks, from healthcare to finance, demonstrate that querying LLMs about the conditional probabilities of events provides meaningful results when compared to baselines, including random and uniform distributions, as well as approaches based on next-token generation probabilities. We explore how these LLM-derived distributions can serve as expert priors to refine distributions extracted from minimal data, significantly reducing systematic biases. Overall, this work introduces a promising strategy for automatically constructing Bayesian Networks by combining probabilistic knowledge extracted from LLMs with small amounts of real-world data. Additionally, we evaluate several prompting strategies for eliciting probabilistic knowledge from LLMs and establish the first comprehensive baseline for assessing LLM performance in extracting probabilistic knowledge.
Extracting Probabilistic Knowledge from Large Language Models for Bayesian Network Parameterization
Aliakbar Nafar1, Kristen Brent Venable2,3, Zijun Cui1, Parisa Kordjamshidi1 1Michigan State University 2Florida Institute for Human and Machine Cognition (IHMC) 3University of West Florida {nafarali, cuizijun, kordjams}@msu.edu, bvenable@ihmc.org
@@ 1 Introduction
The potential of language models as sources for extracting factual knowledge has been demonstrated in several studies (Petroni et al., 2019; Roberts et al., 2020; AlKhamissi et al., 2022; Zhao et al., 2025). However, it remains unclear whether LLMs possess the ability to generate meaningful probabilistic estimates for events and their relationships based on their internal knowledge. We refer to an LLM’s capability to assign a specific probability to an uncertain proposition by utilizing its internal knowledge as probabilistic estimation. This contrasts with providing a confidence score for a concrete answer (Xiong et al., 2024) or solving a probabilistic problem that has a known numerical solution, a process referred to as probabilistic inference (Nafar et al., 2024b). For instance, consider the question, “What is the probability that a person who smokes cigarettes will develop cancer in their lifetime?” There is no definitive correct answer to infer directly from the question itself; however, a medical expert familiar with the literature might approximate 20%. Similarly, we expect an adept language model to produce a similar estimate.
Here, we evaluate the probabilistic estimation capabilities of LLMs such as GPT-4o (OpenAI et al., 2024), Gemini Pro 1.5 (Team et al., 2024), and Claude 3.5 (Anthropic, 2024), and utilize their internal knowledge to construct domain-specific Bayesian networks (BNs). BNs are a powerful method for representing uncertainty and dependencies between events. The reliability of inference in BNs hinges on the accuracy of conditional probability table (CPT) entries, which are typically obtained by collecting data or eliciting expert judgments. These approaches can be expensive or unattainable in domains where data is scarce (You et al., 2019; Longato et al., 2023) or expert time is limited (Das, 2008; Xiaoguang et al., 2019).
By extracting probabilistic knowledge of the LLMs, we explore whether they can serve the role of experts for parameterizing BNs. We assume the dependency structure within the BN is given, leaving the task of extracting such dependencies using LLMs for future research. Nonetheless, preliminary results indicate that LLMs show promise in accurately identifying network dependencies (Babakov et al., 2025). Given the structure of a BN, LLMs are required to predict a probability distribution for each node, conditioned on its parent nodes. In this work, we focus on discrete variables, so we query the LLM separately for each state of a node, prompting it to directly generate a numerical output representing the conditional probability given its parent nodes. After collecting these raw numerical outputs from the LLM, a normalization step is performed to ensure valid probability distributions.
In domains where only a small dataset is available, initial distributions estimated by LLMs can potentially function as expert-derived prior probabilities. We test this hypothesis by adjusting the LLM predictions with a small number of data samples, effectively applying a partial calibration to the model’s initial estimates. We demonstrate that this approach not only enhances the predictions of the LLMs but also surpasses purely data-driven methods and provides a superior prior compared to a uniform prior baseline, which is typically employed in the absence of additional information. Furthermore, our experiments indicate that even when the number of data samples is large, incorporating LLM priors improves the performance. These findings highlight the promise of leveraging LLMs as expert knowledge sources for probabilistic estimation across various real-world domains.
@@
1) Comprehensive Evaluation of LLM Probabilistic Estimation with Real-World BNs. We provide the first large-scale empirical analyses of how LLMs quantify probabilities. We investigate differences in LLM accuracy across diverse BN domains and varying levels of conditional complexity, highlighting the strengths and limitations of LLMs as probabilistic knowledge bases.
2) LLM Predictions as Expert-Driven Priors. We demonstrate that the LLM predicted probabilities can serve as expert-derived priors and combining them with sampled data improves accuracy and sample efficiency compared to purely data-driven methods or using a uniform prior baseline.
[…]
@@ 2 Related Work
Parameter Learning. Within the field of Bayesian Networks, parameter learning has historically revolved around data-driven approaches (Ji et al., 2015) or expert elicitation (Mendes, 2014). Data-driven methods compute frequencies for each configuration of parent states, estimating conditional probabilities via frequentist or Bayesian methods. Expert elicitation is commonly used in domains where data is limited or unobserved for some variables (Woodberry et al., 2005; Constantinou et al., 2016). This work aims to see whether LLM-based probability queries can fill a similar role to experts.
Zero-shot Regression. Using LLMs for regression in a zero-shot setting is an emerging field, and work in this area is limited. Following Vacareanu et al. (2024), which shows that LLMs are capable regressors in a few-shot setting, Nafar et al. (2025) tests the regression capability of LLMs in a zero-shot setting (using internal knowledge) for realistic questions such as estimating the medical insurance cost based on age. However, they don’t use any probability estimation. Paruchuri et al. (2024) asks LLMs to calculate the probabilities for a range of values in a given distribution, but their dataset is limited to only 12 questions, and the queries are elementary with no conditional variables.
@@
Confidence Elicitation. Confidence elicitation in LLMs has been studied in classification tasks, where a confidence score ranging from 0 to 1 is assigned to a discrete class label. Among these, Kadavath et al. (2022) treat the models as a white box and use their token probabilities to assess confidence. However, token likelihood merely indicates the model’s uncertainty about the next token (Kuhn et al., 2023), rather than the confidence in the text itself. Consequently, Xiong et al. (2024); Yang et al. (2024) treat the model as a black-box and use its generated confidence to solve classification datasets. However, confidence elicitation does not directly apply to scenarios requiring a probability distribution across multiple states. Moreover, while similar, confidence is inherently different from probability (Levine, 2024). Another stream of research, in which confidence aligns more closely with probability, prompts LLMs to produce binary probabilities for intermediate nodes, which are also created by LLMs, forming simplified graphical models tailored toward classification tasks (Huang et al., 2025; Feng et al., 2025). In contrast to these methods—dependent on smaller, synthetic graphs constructed explicitly for classification, our approach utilizes large-scale, real-world networks necessitating probability distributions across multiple states. Furthermore, while Feng et al. (2025) extracts hedge words from LLM outputs and subsequently trains a separate model to estimate corresponding probabilities, we directly prompt the LLM to provide explicit probability values.
@@
Probabilistic Inference. Probabilistic inference is closely related to our task and can be considered a natural extension of probabilistic estimation. Saeed et al. (2021); Nafar et al. (2024a) fine-tune BERT-based language models to perform probabilistic inference, while Nafar et al. (2024b) utilizes prompt engineering techniques to enable LLMs to conduct probabilistic inference. However, in all these approaches, the explicit probabilities of sentences are either provided in the text or learned from the dataset during fine-tuning without any estimation derived from the prior internal knowledge.
@@ 3 Problem Definition
The main problem addressed in this paper is parameterizing a Bayesian Network given its structure. We formally define the problem as follows, given the structure of a Bayesian Network $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $V$ is the set of nodes (random variables) and $E$ is the set of edges (dependencies among variables), the goal is to estimate the parameters, that is, Conditional Probability Tables (CPTs) $\mathcal{G}_{\theta}$ of this network using LLMs. We investigate whether LLMs can be used to obtain these parameters. Specifically, the problem becomes one of posing textual queries describing a node’s parents and requesting a probability distribution for each row of the CPT. After this parameter assignment step, we compare the resulting distribution to ground-truth values of the original Bayesian Networks.
@@ 4 Methodology
4.1 Overview
We introduce two distinct frameworks for estimating the parameters of a BN. Our first approach, Extracting Probabilistic Knowledge (EPK), uses prompting techniques to derive probabilistic estimates. Instead of querying the entire probability distribution at once, we individually prompt the LLM for probabilities corresponding to each state of a node and subsequently normalize these values to ensure a valid distribution. Our second approach, Expert-Driven Priors (EDP), assumes the availability of a small empirical dataset and incorporates the LLM-generated predictions as expert-driven priors combined with the empirical data.
4.2 Extracting Probabilistic Knowledge (EPK)
Figure 1: BN Parameterization using EPK. The left side represents an input BN with its dependency structure. The right side shows how EPK uses LLMs to parameterize the BN. Given a node (e.g., Lung Cancer) and its parents, the LLM is prompted to estimate probabilities of its states. The node description is appended to each question to clarify the meanings and states. The raw probabilities generated by the LLM are then normalized to sum to one.
@@ 4.2 Extracting Probabilistic Knowledge (EPK)
As depicted in Figure 1, the process starts with a prompt template that describes the node and its parents in a natural language format and instructs the LLM to provide a probability value. This step is necessary to establish the nodes’ meaning and possible values. The descriptions of these nodes, combined with the LLM instructions, are appended to each question presented to the LLM. Each question explicitly defines the states of the node of interest as well as the states of its parent nodes, and poses a probabilistic query based on these assigned states. Depending on the instruction, the LLM can produce the probability directly or articulate its reasoning before generating a probability value. In both settings, the final answer is extracted as a numerical probability from the output text.
@@
In the questions posed to the LLM in Figure 1, providing descriptions for nodes and their states is optional when the node names and their corresponding values are straightforward. For example, the meanings of “Lung Cancer” and “Smoking” is intuitive, and the LLM can infer that their values are binary, based on the given assignments. However, in cases where the meaning or value sets of nodes are not immediately clear, explicit descriptions are essential. For instance, a node named “X1” must have a clearly stated meaning, such as: “Represents a lack of supervision and policy guidance, which may lead to the use of unqualified oil. This node can take on the values True or False.” Similarly, while the meaning of the “Construction Year” node is self-evident, its possible states are ambiguous. Providing a description like “This node indicates the time period in which the building was constructed, with possible values being 1930-1955, 1955-1960, 1960-1968, 1968-1975, and 1975-1980.”
Since the raw numeric outputs may not sum to one, they are normalized to form a valid distribution over the node’s states. For a node with $m$ states, the model might produce values $p_{1},p_{2},\dots,p_{m}$ that sum to $S=\sum_{i=1}^{m}p_{i}$. To convert these values into a valid probability distribution, we divide each one by $S$. This normalization step can be interpreted as taking the ratio of each state’s assigned likelihood relative to the sum of all states, effectively preserving the proportions while enforcing a valid distribution.
4.3 Expert-Driven Priors (EDP)
Expert opinion and large datasets are either cost-prohibitive or difficult to obtain in many practical scenarios. However, collecting a small number of real observations may still be feasible. When only limited data is available, incorporating expert prior knowledge can help offset the shortage of empirical data, thereby improving the resulting probability distributions. We propose that LLMs can fill the role of experts by providing these prior distributions based on their knowledge. In our approach, EDP, we combine the LLM-derived probabilities with the empirical distribution estimated from these few samples by utilizing two approaches: linear pooling (Gneiting and Ranjan, 2013) and using priors as pseudocounts (Zhai and Lafferty, 2001). For linear pooling, let $q_{1},\dots,q_{m}$ denote the probabilities assigned by the LLM (after normalization) for a particular node with $m$ possible states under a specific parent configuration, and let $r_{1},\dots,r_{m}$ represent the empirical probabilities derived using maximum likelihood estimation, obtained from the collected sample data. To form a final distribution, we apply a linear combination where each final probability is given by $p_{i}=\alpha q_{i}+(1-\alpha)r_{i}$ for $i=1,\dots,m$. Here, $\alpha\in[0,1]$ is a parameter that determines the weight of LLM-derived distribution relative to the empirical one. In the approach where we use the priors as pseudocounts, we translate the LLM-derived probabilities $q_{1},\dots,q_{m}$ into pseudocounts representing prior knowledge by introducing a confidence hyperparameter $C$ and computing pseudocounts $C\cdot q_{i}$ for each state $i$. Then, we combine these pseudocounts with observed counts $n_{i}$ from collected data, obtaining updated probabilities $p_{i}=(C\cdot q_{i}+n_{i})/(C+\sum_{j=1}^{m}n_{j})$ for $i=1,\dots,m$. The hyperparameters $\alpha$ and $C$, which determine the effectiveness of priors, are set to be the reciprocal of data sample numbers, reflecting the intuition that their effect should diminish with more data.

### v2

@@ Abstract
In this work, we evaluate the potential of Large Language Models (LLMs) in building Bayesian Networks (BNs) by approximating domain expert priors. LLMs have demonstrated potential as factual knowledge bases; however, their capability to generate probabilistic knowledge about real-world events remains understudied. We explore utilizing the probabilistic knowledge inherent in LLMs to derive probability estimates for statements regarding events and their relationships within a BN. Using LLMs in this context allows for the parameterization of BNs, enabling probabilistic modeling within specific domains. Our experiments on eighty publicly available Bayesian Networks, from healthcare to finance, demonstrate that querying LLMs about the conditional probabilities of events provides meaningful results when compared to baselines, including random and uniform distributions, as well as approaches based on next-token generation probabilities. We explore how these LLM-derived distributions can serve as expert priors to refine distributions extracted from data, especially when data is scarce. Overall, this work introduces a promising strategy for automatically constructing Bayesian Networks by combining probabilistic knowledge extracted from LLMs with real-world data. Additionally, we establish the first comprehensive baseline for assessing LLM performance in extracting probabilistic knowledge.
@@ 1 Introduction
Bayesian Networks (BNs) are a powerful formalism for representing uncertainty and dependencies between events. The reliability of inference in BNs hinges on the accuracy of the conditional probability table (CPT) entries. CPTs are typically obtained by collecting data (Ji, Xia, and Meng 2015), which can be expensive or unattainable in domains where data is scarce (You, Li, and Shen 2019; Longato et al. 2023). When data is limited, expert judgments are used as priors to be combined with the data for more accurate probability estimation (Mendes 2014). However, experts are often unavailable (Das 2008; Xiaoguang, Yu, and Zhigao 2019); when they are, their competence must be vetted (Hald et al. 2016), and their opinions must be aggregated (McAndrew et al. 2021) before their views can be used. Given these issues, we explore the capability of Large Language Models (LLMs) to act as experts for building BNs.
@@
The potential of language models as sources for extracting factual knowledge has been demonstrated in several studies (Petroni et al. 2019; Roberts, Raffel, and Shazeer 2020; AlKhamissi et al. 2022; Zhao et al. 2025). However, it remains unclear whether LLMs possess the ability to generate meaningful probabilistic estimates for events and their relationships based on their internal knowledge. In this paper, we use the term probabilistic estimation to refer to assigning a specific probability to an uncertain proposition by LLM when utilizing its internal knowledge. For instance, consider the question, “What is the probability that a person who smokes cigarettes will develop cancer in their lifetime?” The answer to this question cannot be inferred from the question’s context; however, a medical expert familiar with the literature might approximate 20%. Similarly, we expect an adept language model to produce a similar estimate. This contrasts with providing a confidence score for a concrete answer (Xiong et al. 2024) or solving a problem that has a known numerical solution.
We evaluate the probabilistic estimation capabilities of LLMs such as GPT-4o (OpenAI et al. 2024), Gemini Pro 1.5 (Gemini Team et al. 2024), Claude 3.5 (Anthropic 2024), and open-source model DeepSeek-V3 (DeepSeek-AI et al. 2025b) and utilize their internal knowledge to construct domain-specific BNs with discrete variables. To provide a detailed analysis and clear evaluation of the parameter estimation, we assume the dependency structure within the BN is given. We note that the dependency structure extraction has been investigated in (Babakov, Reiter, and Bugarín-Diz 2025) and the preliminary results are promising. In our setting, given the structure of a BN, LLMs are required to predict a probability distribution for each node, conditioned on its parent nodes. First, we analyze the quality of the initial distributions estimated by LLMs. Then, we investigate whether they can potentially function as expert-derived prior probabilities. We test this approach, which we denote as Expert-Driven Priors (EDP), by adjusting the LLM predictions with data samples, effectively applying a partial calibration to the model’s initial estimates.
We use LLMs to estimate the CPTs of 80 real-world BNs collected from the literature. Using Kullback–Leibler (KL) divergence (Kullback and Leibler 1951) as our metric, we show that EDP consistently improves LLM’s predictions and offers a higher‑quality prior than the conventional uniform baseline, which is employed when no extra information is available. Furthermore, our experiments indicate that even when the number of data samples is large, incorporating LLM priors improves performance. We further evaluate EDP on BNs used for classification, demonstrating that reducing KL divergence translates into higher downstream classification accuracy. These findings highlight the promise of leveraging LLMs as expert knowledge sources for probabilistic estimation across various real-world domains.
@@
1) We introduce the first large-scale and comprehensive evaluation of LLM probabilistic estimation with real-world BNs. We investigate differences in LLM accuracy across domains and varying levels of network complexity, highlighting their effectiveness as probabilistic knowledge bases.
2) We show that the LLM predicted probabilities can serve as expert-derived priors, and combining them with data improves the estimated probabilities compared to purely data-driven methods or using a uniform prior baseline.
3) We evaluate our method on downstream classification tasks and demonstrate that the improved probability distributions lead to higher classification accuracy.
4) We introduce an automated procedure for parameterizing real-world BNs given the network structure  [^1] The code and analytical tools are available at https://github.com/HLR/llm-bn-parameterization..
@@ 2 Related Work
To the best of our knowledge, this is the first study to use LLMs to parameterize Bayesian networks and to evaluate those parameters against ground-truth probabilities. Most prior work queries LLMs for a single number interpreted as a “confidence” in a class label, which is inherently different from probabilities (Levine 2024). Among the few works that elicit probabilities, evaluation is only done at task accuracy because gold probabilities are unavailable. In contrast, we directly assess the quality of the learned CPTs via KL divergence to the original BN parameters, and in addition, examine downstream classification accuracy.
Probability Estimation. In the most relevant prior work, LLMs are prompted to produce probabilities of binary variables forming shallow BNs with a depth of only two, tailored toward classification tasks (Huang et al. 2025; Feng et al. 2025). Although they show improved downstream classification accuracy, they do not evaluate the full probability distributions because gold probabilities are unavailable. Feng et al. (2025) only evaluates the correctness of the magnitudes of inferred-node probabilities, not having the ground-truth as a basis to evaluate the distributions. By contrast, we obtain a complete probability distribution for every discrete node, including those with more than two states, and evaluate it against real‑world BNs of varying depth with known parameters. Further, we show how an improved estimation translates into better downstream classification Tasks. Paruchuri et al. (2024) asks LLMs to calculate the probabilities for a range of values in a given distribution, but their dataset is limited to only 12 questions, and the queries are elementary with no conditional probabilities.
@@
Confidence Elicitation. Confidence elicitation in LLMs has been studied in classification tasks, where a confidence score ranging from 0 to 1 is assigned to a discrete class label. Among these, Kadavath et al. (2022) treat the models as a white box and use their token probabilities to assess the confidence of a label. But, token likelihood indicates the model’s uncertainty about the next token (Kuhn, Gal, and Farquhar 2023), rather than the confidence of the label. Consequently, Xiong et al. (2024); Yang, Tsai, and Yamada (2024) treat the model as a black-box and use its generated confidence to solve classification datasets. However, confidence is different from probability (Levine 2024) and confidence elicitation does not directly apply to scenarios requiring a probability distribution across multiple states.
@@
Probabilistic Inference. Probabilistic inference is closely related to our task and can be considered a natural extension of probabilistic estimation. Saeed et al. (2021); Nafar, Venable, and Kordjamshidi (2024b) fine-tune BERT-based language models to perform probabilistic inference, while Nafar, Venable, and Kordjamshidi (2024a) utilizes prompt engineering techniques to enable LLMs to conduct probabilistic inference. However, in all these approaches, the explicit probabilities are either provided in the text or learned from the dataset during fine-tuning without any estimation derived from the internal knowledge of language models.
Zero-shot Regression. Our method queries the LLMs for a numeric probability expressed in plain text, in a relatively similar setting to using LLMs for regression. Using LLMs for regression in a zero-shot setting is an emerging field, with a limited number of studies. Following (Vacareanu et al. 2024), which shows that LLMs are capable regressors in a few-shot setting, Nafar, Venable, and Kordjamshidi (2025) tests the regression capability of LLMs in a zero-shot setting (using internal knowledge) for realistic questions such as estimating the medical insurance cost based on age. However, they don’t use any probability estimation.
@@ 3 Problem Definition
The main problem addressed in this paper is parameterizing a Bayesian Network given its structure. We formally define the problem as follows, given the structure of a BN $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $V$ is the set of nodes (random variables) and $E$ is the set of edges (dependencies among variables), the goal is to estimate the parameters, that is, Conditional Probability Tables $\mathcal{G}_{\theta}$ of this network with the help of LLMs. After assigning the parameters, we compare the resulting distribution to ground-truth values of the original BNs.
@@ 4 Methodology
Figure 1: Two‑stage parameterization pipeline. SepState (Left Panel): For each parent configuration, the LLM is prompted with natural‑language descriptions of the node and its parents and queried once per state. The answers are subsequently normalized ($z^{-1}$) into a valid conditional distribution, e.g., (75%, 25%). EDP (Right Panel): The LLM‑derived prior distribution is translated into pseudocounts and fused with empirical counts ($C$) to give the posterior estimates. This treats the LLM as a probabilistic expert whose influence is controlled by the hyper‑parameter $\alpha$.
As illustrated in Figure 1, our BN-parameterization pipeline has two steps. First, we query an LLM to use its internal probabilistic knowledge to generate probability estimates. Next, we use these estimates in our Expert-Driven Priors (EDP) framework, which treats the LLM outputs as priors and refines them with data to produce the final parameters.
Extracting Probabilistic Knowledge
@@
In the first stage of our BN-parameterization pipeline, we use LLMs to acquire probabilistic estimates for every row of the CPTs. Because many nodes are multi-state, asking for a single probability score for each node is impractical. So we use two prompting schemes: 1) FullDist, where the entire distribution is obtained from the LLM at once in the form of a tuple, e.g., $(0.70,0.20,0.10)$. 2) SepState, where each state is queried independently. Figure 1 depicts the SepState scheme, where the process starts with a prompt template that describes the node and its parents in a natural language format. The descriptions of these nodes, combined with the LLM instructions, are appended to each question presented to the LLM. Each question explicitly defines the states of the node of interest as well as the states of its parent nodes, and poses a probabilistic query based on these assigned states. The LLM is also instructed to articulate its reasoning before generating a probability value. The final answer is extracted as a numerical probability from the output text. Since the raw numeric outputs may not sum to one, they are normalized to form a valid distribution over the node’s states. For a node with $m$ states, the model might produce values $p_{1},p_{2},\dots,p_{m}$ that sum to $S=\sum_{i=1}^{m}p_{i}$. To convert these values into a valid probability distribution, we divide each one by $S$. This normalization step can be interpreted as taking the ratio of each state’s assigned likelihood relative to the sum of all states, effectively preserving the proportions while enforcing a valid distribution.
FullDist follows the same prompting template, but the LLM is instructed to return the entire distribution in one shot. For example, instead of the two questions in Figure 1, the LLM is asked “What is the probability distribution of lung cancer given that smoking is true?” Thus, the LLM returns a tuple $(p_{1},\dots,p_{m})$ in a single response. Unlike SepState, the FullDist scheme is concise and requires only one query, independent of the number of the node’s states. However, the autoregressive decoder conditions each number on the previous ones, so early outputs can bias later entries. Concentrating all states in a single prompt could also be too complicated for the LLM. In our experiments, we compare SepState and FullDist schemes empirically.
@@
As noted above, our prompts include brief definitions of each node and its state space. However, this information can sometimes be extracted from the questions themselves. For example, the meanings of “Lung Cancer” and “Smoking” are commonsense knowledge and known to the LLMs. Also, the LLM can infer that their values are binary, based on the given true/false assignments. However, in cases where the semantic meaning or value sets of nodes are not immediately clear, explicit descriptions are needed. For instance, a node named “X1” must have a clearly stated meaning, such as: “Represents a lack of supervision and policy guidance, which may lead to the use of unqualified oil. This node can take True or False values.” Similarly, while the semantic meaning of the “Construction Year” node is self-descriptive, its possible states are ambiguous and need an explanation such as: “This node indicates the time period in which the building was constructed, with possible values being 1930-1955, 1955-1960, 1960-1968, 1968-1975, and 1975-1980.” We later test the usage of the contextual descriptions of node meanings and state sets by LLMs in an ablation study to measure their effect on the LLM’s probability predictions.
Expert-Driven Priors (EDP)
Expert opinion and large datasets are either costly or difficult to obtain in many practical scenarios. When only limited data is available, incorporating expert prior knowledge can particularly help offset the shortage of empirical data, thereby improving the estimated probability distributions. We propose that LLMs can approximate these prior distributions based on their knowledge. In our approach, EDP, we combine the LLM-derived probabilities with the empirical distribution estimated from data by using priors as pseudocounts (Zhai and Lafferty 2001).
For a node in the BN with $m$ discrete states, conditioned on a specific parent configuration, the LLM provides a normalized prior distribution $q_{1},\dots,q_{m}$ as shown in the left panel of Figure 1. Then, we use these priors and incorporate data to calculate the final probability distribution. Let $c_{1},\dots,c_{m}$ denote the observed counts listed in the data table on the right side of Figure 1. Each prior probability $q_{i}$ is converted into $\alpha q_{i}$ virtual observations, where $\alpha$ is a hyperparameter. Larger $\alpha$ values place more weight on the LLM prior, whereas smaller values defer to the data. Adding these pseudocounts to the empirical counts gives the posterior estimate displayed inside the dashed box of Figure 1:
$p_{i}=\frac{\alpha q_{i}+c_{i}}{\alpha+\sum_{j=1}^{m}c_{j}},\qquad i=1,\dots,m.$
The numerator combines prior belief $(\alpha q_{i})$ with empirical evidence $(c_{i})$, while the denominator $\alpha+\sum_{j}c_{j}$ normalizes the resulting probabilities. In this manner, we incorporate LLM predictions in the same principled way one would incorporate probabilities elicited from a human expert.